OpenStudy (anonymous):

the derivative of square root of 6x is 3x^(-1/2)?

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

do you need explanation?

OpenStudy (anonymous):

i got it but its telling me its not correct? any ideas why?

OpenStudy (turingtest):

derivative of\[\sqrt{6x}=(1/2)[(6x)^(-1/2)](6)=3[(6x)^(-1/2)]=3/\sqrt{6x}\] oh I guess it was wrong...